# Automated Testing Method for Text-to-Image Software

### v2

@@ Abstract.
Recently, creative generative artificial intelligence software has emerged as a pivotal assistant, enabling users to generate content and seek inspiration rapidly. Text-to-image (T2I) software, being one of the most widely used among them, is used to synthesize images with simple text input by engaging in a cross-modal process. However, T2I software often encounters defects and erroneous, including omitting focal entities, low image realism, and mismatched text-image information. The cross-modal nature of T2I software makes it challenging for testing methods to detect defects. Lacking test oracles further increases the complexity of testing.
To address this deficiency, we propose ACTesting, an Automated Cross-modal Testing Method of Text-to-Image software, the first testing method designed specifically for T2I software. We construct test samples based on entities and relationship triples following the fundamental principle of maintaining consistency in the semantic information to overcome the cross-modal matching challenges. To address the issue of testing oracle scarcity, we first design the metamorphic relation for T2I software and implement three types of mutation operators guided by adaptability density. In the experiment, we conduct ACTesting on four widely-used T2I software. The results show that ACTesting can generate error-revealing tests, reducing the text-image consistency by up to 20% compared with the baseline. We also conduct the ablation study that effectively showcases the efficacy of each mutation operator based on the proposed metamorphic relation.
@@
ccs: Software and its engineering Software testing and debugging
@@ 1. Introduction
Deep learning has undergone significant evolution. The rise in popularity of transformer (Devlin et al., 2018), generative adversarial network (GAN) (Zhu et al., 2017) and diffusion model (Rombach et al., 2022) empower Artificial Intelligence Generative Content (AIGC) to yield surprising results in fields such as image, text, and audio generation within simple inputs. The emergence of large-scale models precipitates a rise in creative generative tasks. Text-to-image (T2I) task, being one of the most popular among them, aims to automatically synthesize a creative image based on the simple input text as shown in Figure 1. Recently, several T2I models (Zhang et al., 2017; Huang et al., 2019; Lao et al., 2019; Gu et al., 2022; Saharia et al., 2022; Qiao et al., 2019; Rombach et al., 2022; Kim et al., 2022) and software (Baidu, 2023; Alibaba, 2023; OpenAI, 2023; Stability.ai, 2023; DeepAI, 2023; Midjourney, 2023) from top IT companies show great performance on fidelity and creativity of the output image. The high-quality content and the simple interaction flow make increasing use of T2I software in our daily lives. Representative applications include generating the image for cross-modal data augmentation (Trabucco et al., 2023) (Computer Science), artistic creation and designing (Wu, 2022) (Art) and multilingual communication assistant for deaf (Huynh et al., 2022) (Medical). The ability to handle cross-modal information and understand rich scenarios greatly enhances T2I software’s potential for development.
@@
However, the outputs from T2I software are not entirely reliable or capable of meeting the expected requirements. Despite diverse techniques that have been researched and adapted to improve the inner engine of T2I software, the generative results could be abnormal or incorrect (Hinz et al., 2020; Wu et al., 2022). This is due to the huge and complex neural network structure, labeling errors in training datasets and uneven feature distribution. Generating unpleasant, offensive or inappropriate image content may result in significant repercussions (Tredinnick and Laybats, 2023). Objectionable or discriminatory content can lead to brand damage in a business setting and even trigger public protests or legal action. Consequently, it is significant to test the generation quality and robustness of T2I software.
@@
Effective testing methods can swiftly detect the defects in T2I software, helping measure its robustness. However, developing testing methods for T2I software presents persistent challenges (Frolov et al., 2021; Borji, 2019; Theis et al., 2015). Initially, in contrast to conventional software with delineated internal logic, a significant segment of commercial T2I software offers only API interfaces to end-users, obscuring their internal engine and foundational models. This makes white-box testing methods no longer suitable for this task. Secondly, the most intricate issue is cross-modal alignment, necessitating aligning two modes with completely different information representations during the T2I software’s testing phase. Traditional robustness testing methods, including boundary value analysis, adversarial attack, and anomaly detection, are also not suitable for the cross-modal generation task because of the lack of testing oracles. In addition, classic test methods based on text augmentation such as swapping the order of the words or introducing spelling errors are possible to cause errors in semantic information or affect the quality of the output images. Lastly, the tasks of quantifying the software’s generation robustness and defining anomalies in the testing process are both crucial and demanding. To the best of our knowledge, there are no testing methods specifically designed for T2I software available.
To tackle the challenges mentioned above, we propose ACTesting, an automated and black-box cross-modal testing method for T2I software. ACTesting is designed based on the metamorphic testing theory to address the issue of oracle scarcity and cross-modal semantic matching. Specifically, it designs the metamorphic relation based on the fundamental principle of maintaining consistency in the semantic information across different modalities. To keep the semantic preservation, we apply the entity-relationship (ER) triple to represent the focal semantic information in both modalities. Then, we implement three types of mutation operators based on the ER triple to detect the defects of tested T2I software. This method aims to test the generation robustness of such software for consistently high quality and cross-modal alignment output.
@@
We conduct experiments on four widely-used T2I software to evaluate the effectiveness of ACTesting. Tested software generates 113,736 synthetic images based on the MS-COCO validation dataset referring (Dinh et al., 2022). We employ I-FID, I-IS and RP ((Bynagari, 2019; Salimans et al., 2016; Xu et al., 2018; Dinh et al., 2022)) as the basic evaluation metrics for image realism and text-image relevance. We also design the error-detecting ($Error_{e}$ and $Error_{r}$) and miss-detection ($Miss_{e}$ and $Miss_{r}$) rates as the metrics, based on the designed metamorphic relation (Yu et al., 2022b) and scene graph generation model (Tang et al., 2020). The experimental results demonstrate that ACTesting can reduce image quality by 2.9% to 15% and decrease text-image match consistency by 7.5% to 21.1%. The average error rate of the three operators is around 60%, which is 1.75 times higher than that of the baseline text mutation operator. Moreover, we conduct the ablation study to further elucidate the effectiveness of each operator. Not only can the operators be flexibly combined, but the results also show that these combined operators effectively increase the miss-detection rate beyond that of the basic operators.
The main contributions of this paper are as follows:
@@
  • Method. We introduce the first automated cross-modal testing method for T2I software, termed ACTesting. This method keeps the core idea of cross-modal semantic preservation and employs the entity-relationship triple to design the metamorphic relation. We design three kinds of mutation operators to detect the erroneous and test the generation robustness of T2I software under the guidance of adaptability density.
@@
  • Tool. We integrate the aforementioned method, ACTesting, into a Python tool. This represents the first black-box automated testing tool for T2I. We make the code for the entire process available on GitHub [^1] https://anonymous.4open.science/r/ACTesting-9478/.
@@
  • Study. We conduct a comprehensive experiment to evaluate the performance of four industrial T2I software utilizing our ACTesting. The test findings further illustrate that our testing procedure can adeptly produce error-revealing test cases, leveraging our adaptability density-guided operators. The ablation study further shows that the proposed mutation operator can be flexibly combined for enhanced effectiveness.
2. Preliminaries
@@
In this section, we begin by outlining the workflow of modern Text-to-Image (T2I) software. We then detail the techniques employed by current T2I software, introducing how they effectively convert text into images. Finally, we present the motivation behind our proposal. This introduction sets the stage for a deeper exploration of our innovative contributions to testing T2I software.
@@ 2.1. T2I Software Workflow
Figure 2. The workflow of modern T2I software
@@
Figure 2 delineates the general pipeline of modern T2I software. Leveraging expansive training image-text datasets, text encoder and image encoder models undergo pre-training to extract the cross-modal feature and establish mappings within a latent space. After that, the image generator and decoder models are trained end-to-end for image information reconstitution. T2I software then deploys these trained models (the text encoder, image generator, and image decoder) for the image synthesis procedure. For more flexible adjustment of input and output, T2I software designs pre-processing and post-processing stages to verify data quality. The pre-processing stage not only ensures text input conforms to predefined standards but also augments more details to the input text to fit the trained models. The initial image is generated based on the processed text input. In the terminal phase, post-processing assesses the initial image for both fidelity and ethical considerations, further employing image enhancement algorithms to mitigate noise and fulfill details. The core part of the whole workflow is the trained T2I model, thus we briefly introduce some modern engines in the next section.
2.2. Modern T2I Engine
@@
With the rapid development of this technique, we introduce two main series of popular T2I models: Autoregressive and Diffusion-based models.
@@
Autoregressive Model. Autoregressive methods can exploit large-scale datasets through time-series models, predicated on the dependence of the historical time-series of forecasting targets in different periods. Representative method DALL-E (Ramesh et al., 2021) from OpenAI, which regards both text and image tokens as sequential data, performing autoregression via Transformer architecture. Another model Parti (Yu et al., 2022a) from Google encodes images as sequences of discrete tokens and reconstructs these sequences into high-quality images. However autoregressive methods nature need substantial computation resources.
@@
Diffusion-Based Model. The most predominant approach is diffusion-based methods, which are also the new state-of-the-art models in T2I generation. Diffusion models (DMs) aim to reserve a process of perturbing the data with noise for sample generation. As a milestone work, Stable diffusion (Rombach et al., 2022) trains the diffusion models within latent space, incorporating the text tokens during the denoising phase to fuse the cross-model features. DALL-E2 (Ramesh et al., 2022) applies CLIP (Radford et al., 2021) to learn and match the multimodal representations, evidencing unparalleled efficacy. Predominantly, contemporary T2I software engines are grounded in diffusion models.
@@ 2.3. Motivation
Figure 3. Examples of the input-output pairs of text-to-image software. All the images are generated with OpenAI.DALL-E
As depicted in Figure 3, we present examples generated by prevalent T2I software. In the left part (Figure a), we employ four variations of the sentence as inputs. Although the input text conveys nearly identical meanings, the output images exhibit variations in missing components. Subfigures 1 and 3 omit the ”man”, while subfigure 2 lacks the ”bird”. Only subfigure 4 accurately represents the correct entities (”man”, ”bird”, ”stone”) and relationships (”watching”, ”standing on”). Moreover, in the right part (Figure b), we note that with an increase in the number of entities and relationships, the realism of the images diminishes. There are further omissions evident in both entities (subfigure 5) and relationships (subfigure 6).
Therefore, we consider that T2I software exhibits high sensitivity to the focal ERs, where even simple alterations in sentence structure could potentially lead to a reduction in text-image consistency. This implies that the design of metamorphic relations based on cross-modal semantic preservation should also be approached from the perspectives of entities and relationships.
@@ 3. Approach
Figure 4. The Overview of ACTesting
This section introduces ACTesting, which is proposed to test T2I software automatically and evaluate the generation robustness of the software in a black-box scenario. Figure 4 presents the overview of ACTesting. ACTesting can generate new input text based on the seed text by applying transformation operators. Rather than utilizing common text augmentation methods, ACTesting implements adaptability density constraint to design the operators of replacing, removing, and augmenting the ER triple. ACTesting designs metamorphic relation to testing the generation robustness based on the image entity and space relationship detection method.
3.1. Adaptability Density-guided Mutation Operators

### v3

# ACTesting: Automated Cross-modal Testing Method of Text-to-Image Software

@@ Abstract.
Recently, creative generative artificial intelligence software has emerged as a pivotal assistant, enabling users to generate content and seek inspiration rapidly. Text-to-Image (T2I) software, one of the most widely used, synthesizes images with text input by engaging in a cross-modal process. However, despite substantial advancements in the T2I engine, T2I software still encounters errors when generating complex or non-realistic scenes, including omitting focal entities, low image realism, and mismatched text-image information. The cross-modal nature of T2I software complicates error detection for traditional testing methods, and the absence of test oracles further exacerbates the complexity of the testing process. To fill this gap, we propose ACTesting, an Automated Cross-modal Testing Method of Text-to-Image Software, the first testing method explicitly designed for T2I software. ACTesting utilizes the metamorphic testing principle to address the oracle problem and identifies cross-modal semantic consistency as its fundamental Metamorphic relation (MR) by employing the Entity-relationship (ER) triples. We design three kinds of mutation operators under the guidance of MR and the adaptability density constraint to construct the new input text. After generating the images based on the text, ACTesting verifies whether MR is satisfied by detecting the ER triples across two modalities to detect the errors of T2I software. In our experiments across five popular T2I software, ACTesting effectively generates error-revealing tests, resulting in a decrease in text-image consistency by up to 20% when compared to the baseline. Additionally, an ablation study demonstrates the efficacy of the proposed mutation operators. The experimental results validate that ACTesting can reliably identify errors within T2I software.
@@
copyright: acmcopyrightjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: TOSEMjournalvolume: 0journalnumber: 0article: 1publicationmonth: 0ccs: Software and its engineering Software testing and debugging
@@ 1. Introduction
Deep learning has undergone significant evolution. The rise in popularity of transformer (Devlin et al., 2019), generative adversarial network (GAN) (Zhu et al., 2017) and diffusion model (Rombach et al., 2022) empower Artificial Intelligence Generative Content (AIGC) to yield surprising results in fields, such as image, text, and audio generation within simple inputs. The emergence of large-scale models precipitates a rise in creative generative tasks. Among them, Text-to-Image (T2I) task is one of the most popular ones, aiming to automatically synthesize a creative image based on the simple input text as shown in Fig. 1. Recently, several T2I models (Zhang et al., 2017; Huang et al., 2019; Lao et al., 2019; Gu et al., 2022; Saharia et al., 2022; Qiao et al., 2019; Rombach et al., 2022; Kim et al., 2022) and software (Baidu, 2023; Alibaba, 2023; OpenAI, 2023; Stability.ai, 2023; DeepAI, 2023; Midjourney, 2023) from top IT companies show great performance on fidelity and creativity of the output image. The high-quality content and the simple interaction flow make increasing use of T2I software in our daily lives. Representative applications include generating the image for cross-modal data augmentation (Trabucco et al., 2024) (Computer Science), artistic creation and designing (Wu, 2022) (Art), and multilingual communication assistant for deaf (Huynh et al., 2022) (Medical). The ability to handle cross-modal information and understand rich scenarios greatly enhances T2I software’s potential for development.
@@
However, the outputs from T2I software are not entirely reliable or capable of meeting the expected requirements. Despite diverse techniques that have been researched and adapted to improve the inner engine of T2I software, the generative results could be abnormal or incorrect (Hinz et al., 2020; Wu et al., 2022), especially under unreal-world scenarios. The reasons are the huge and complex neural network structure, labeling errors in training datasets, and uneven feature distribution. Especially, the existing evaluation metrics are unexplainable (mainly based on deep feature similarity) and cannot accurately pinpoint the errors in the generated images. Generating unpleasant, offensive, or inappropriate image content may result in significant repercussions (Tredinnick and Laybats, 2023). Objectionable or discriminatory content can damage a brand in a business setting and even trigger public protests or legal action. Consequently, it is significant to test the generation quality of T2I software.
Figure 1. Sample inputs and outputs of T2I software.
@@
To the best of our knowledge, however, there is no testing method specifically designed for T2I software available. Developing testing methods for T2I software presents persistent challenges  (Frolov et al., 2021; Borji, 2019; Theis et al., 2016). Initially, in contrast to conventional software with delineated internal logic, commercial T2I software kernels are more complex, including both deep learning engines and software modules. Most of the end-users can only get results through the API. This makes white-box testing methods no longer suitable for this task. Secondly, the most intricate issue is cross-modal alignment, necessitating aligning two modes with completely different information representations during the T2I software’s testing phase. Thirdly, the simple testing methods based on text augmentation, such as swapping the order of the words or introducing spelling errors, possibly cause errors in semantic information before generation. In addition, current research (Lee et al., 2024) focuses on quantifying aspects such as image realism or alignment through deep learning models, but they do not explicitly identify defects within the images. Therefore, testers cannot precisely determine which semantic representations in the generated images are problematic.
To tackle the above-mentioned challenges, we propose ACTesting, an automated and black-box cross-modal testing method for T2I software. ACTesting is designed based on the Metamorphic Testing (MT) theory to address the issue of oracle scarcity. The core design concept of Metamorphic relation (MR) in ACTesting is the cross-modal semantic information matching consistency. That is, the semantic information changes of the input text and the output image should be consistent. To construct the MR, we employ the Entity-relationship (ER) model. Specifically, to preserve semantic preservation, we apply the ER triple to represent the focal semantic information in both modalities different from traditional text augmentation. We also propose three mutation operators based on the adaptability density constraint instead of the probability-proportional sampling to construct the new input text. Finally, ACTesting constructs input text based on the proposed mutation operators and treats the MR as practical adequacy criteria to reveal hidden errors of tested T2I software. During the implementation process, ACTesting uses Object Detection and Scene Graph models to estimate the entities and relationships in the generated images, thus eliminating the need for pre-labeling.
@@
To evaluate the effectiveness of ACTesting, we conduct experiments on five widely used T2I software. ACTesting generates 142,170 synthetic images based on the MS-COCO validation dataset (Dinh et al., 2022) in total to test the performance of T2I software. We employ Improved-Fréchet Inception Distance (I-FID), Improved-Inception Score (I-IS), and R-Precision (RP) (Bynagari, 2019; Salimans et al., 2016; Xu et al., 2018; Dinh et al., 2022) as the basic evaluation metrics for image realism and text-image relevance. We also design the error detection ($Error_{e}$ and $Error_{r}$) and miss-detection ($Miss_{e}$ and $Miss_{r}$) rates as the metrics, based on the designed MR  (Yu et al., 2022b) and the scene graph generation model (Tang et al., 2020) to further identify the specific defects within the images. The experimental results demonstrate that ACTesting can improve the detection performance of errors. ACTesting degrades image quality by 2.9% to 15% according to the I-IS metric and reduces text-image match consistency by 7.5% to 21.1% according to the RP metric. The average error rate of the three mutation operators is around 60%, which is 1.75 times higher than the baseline. These results show that ACTesting not only constructs the mutation text samples to reduce the generation quality of the software but also detects specific errors in the generated images. Moreover, we conduct an ablation study to further elucidate the effectiveness of each mutation operator. The results show that the combined mutation operators effectively increase the miss-detection rate beyond that of the single mutation operator. The main contributions of this paper are as follows:
@@
  • Method. We introduce the first automated cross-modal testing method for T2I software based on MT, termed ACTesting. It keeps the core idea of cross-modal semantic preservation and employs the ER triple to construct the MR. We design three kinds of mutation operators under the guidance of adaptability density constraint to detect the errors in the generated images and test the generation quality of T2I software.
@@
  • Tool. We integrate ACTesting into a Python tool. It represents the first black-box automated testing tool for T2I. We make the code and test data of the entire process publicly available at https://github.com/sikygu/ACTesting.
@@
  • Study. We conduct a comprehensive experiment to evaluate the performance of ACTesting on five industrial T2I software. The results illustrate that our testing method can adeptly produce error-revealing test cases, leveraging our adaptability density-guided operators. The ablation study further shows that the proposed mutation operators can be flexibly combined for enhanced effectiveness.
2. Preliminaries and Motivation
@@
In this section, we begin by outlining the workflow of modern T2I software. We then detail the T2I models employed by current T2I software, introducing how they effectively convert text into images. Finally, we present the motivation behind our proposal, which sets the stage for a deeper exploration of our innovative contributions to testing T2I software.
@@ 2.1. T2I Software Workflow
Figure 2. The workflow of modern T2I software.
@@
Fig. 2 delineates the general pipeline of modern T2I software, including training T2I models, generating images, pre-processing, and post-processing. Leveraging expansive training image-text datasets, text encoder and image encoder models undergo pre-training to extract the cross-modal feature and establish mappings within a latent space. After that, the image generator and decoder models are trained end-to-end for image information reconstitution. T2I software then deploys these trained models (the text encoder, image generator, and image decoder) for the image synthesis procedure. For more flexible adjustment of input and output, T2I software designs pre-processing and post-processing stages to verify data quality. The pre-processing stage not only ensures text input conforms to predefined standards but also augments more details to the input text to fit the trained models. The initial image is generated based on the processed text input. In the terminal phase, post-processing assesses the initial image for both fidelity and ethical considerations, further employing image enhancement algorithms to mitigate noise and fulfill details. The core part of the whole workflow is the trained T2I models. Thus, we briefly introduce some modern engines in the next section.
2.2. Modern T2I Models
@@
With the rapid development of T2I models, two main series of T2I models are proposed: Autoregressive and Diffusion-based T2I models. We introduce cutting-edge methods respectively.
@@
Autoregressive Models. Autoregressive methods can exploit large-scale datasets through time-series models predicated on the dependence of the historical time-series of forecasting targets in different periods. Representative method DALL-E (Ramesh et al., 2021) based on CLIP (Radford et al., 2021) from OpenAI regards both text and image tokens as sequential data, performing autoregression via Transformer architecture. Another model, Parti (Yu et al., 2022a) from Google, encodes images as sequences of discrete tokens and reconstructs these sequences into high-quality images.
@@
Diffusion-Based Models. The most predominant approach is diffusion-based methods, which are also the new state-of-the-art models in T2I generation. Diffusion models (DMs) aim to reserve a process of perturbing the data with noise for sample generation. As a milestone work, Stable diffusion (Rombach et al., 2022) trains the diffusion models within latent space, incorporating the text tokens during the denoising phase to fuse the cross-modal features. Predicated Diffusion (Sueyoshi and Matsubara, 2024) is proposed recently, a unified framework designed to more effectively express users’ intentions. Predominantly, contemporary T2I software engines are grounded in diffusion models.
In the experiments, we conduct a comprehensive evaluation by testing five different text-to-image (T2I) software programs. These programs are carefully selected to represent the two distinct types of engines that are widely used in the field. Through this approach, we aim to demonstrate the versatility and broad applicability of ACTesting across diverse T2I generation techniques. By including multiple software platforms with varying architectures and capabilities, we ensure a robust examination of our method’s effectiveness and its potential to adapt to different underlying technologies.
@@ 2.3. Motivation
Figure 3. Examples of the input-output pairs of text-to-image software. The images are generated with OpenAI.DALLE and Stable Diffusion XL.
The motivation of ACTesting is described from the perspective of cross-modal semantics. Fig. 3 shows images generated by prevalent T2I software highlighting the entities in red and the relationships in green. In Fig. 3 (a), two sentence variations with the same semantics are as inputs. Although the input text conveys nearly identical meanings, Image 1 omits the “man” while Image 2 includes the correct entities and relationships. In Fig. 3 (b), the common collocation in the sentence “woman sitting on the bench” is replaced with “woman sitting on the plate”. The generated images demonstrate that collocating unreal-world entity relationships degrades the quality of the generated entity (distortion of hand depiction in Image 4). Moreover, in Fig. 3 (c), complex entities and relationships diminish the realism of the images.
In conclusion, T2I software exhibits high sensitivity to the focal ERs, where even simple alterations in sentence structure could lead to decreased text-image consistency and image quality. This stimulates us to consider that the cross-modal semantic consistency can be used as the core MR of MT in ACTesting. As a result, ACTesting introduces the ER model and using ER triples to be the basic unit of MR.
@@ 3. Approach
Figure 4. The overview of ACTesting.
This section introduces ACTesting, which is proposed to test T2I software automatically and detect the errors of the software in a black-box scenario. ACTesting is based on MT and uses cross-modal semantic consistency as the basic MR. Sec 3.1 introduces the workflow of ACTesting according to Fig. 4. The core concept of ACTesting lies in the adaptability density constraint of T2I software, which quantifies its ability to handle diverse test samples and serves as a measure of generation quality (introduced in Sec 3.2). Then, we characterize the adaptability density constraint by the ER model and define the basic MR in Sec 3.3. To construct the new input text, we design three mutation operators based on the ER model in Sec 3.4.
3.1. Overview of ACTesting
ACTesting is used to construct a complete testing process. As shown in Fig. 4, ACTesting includes several main processing components: ER pool construction, adaptability density-guided operators, text synthesizing, image entity and relationship detection, and testing and evaluation results calculation. We select seed texts from the dataset that contain diverse ERs (according to the class label of the dataset) at the beginning. Then, ACTesting constructs an ER pool for each seed text based on the principle of semantic information preservation.
ER Pool Construction. The input text in the T2I software is different from other natural language processing tasks like text classification or text abstract. The text often contains the scene or the objects of the generated image, which renders traditional text mutation operators (such as swapping the order of the words or introducing spelling errors in the text) no longer applicable to inputs for this task. Considering the different ways various modalities express information, we transform each input text into an ER pool, utilizing consistent cross-modal semantic information features. Subsequently, all mutation operations are performed based on the ER pool, ensuring that the semantic features of the input text remain intact. We adopt the relationship extraction, entity recognition and structured representation technique to construct the ER pool referring (Qi et al., 2020). This method also constructs the ER candidates based on the open-source dataset.
Mutation Operator Implementation. To construct the new testing samples and reserve the cross-modal semantic information, three kinds of mutation operators are designed under the adaptability density-guided constraint (The details of the adaptability density-guided constraint and mutation operators are introduced in Sec 3.2 and Sec 3.4, respectively). ACTesting adopts the EC and ER_A operators to all the inputs and adopts the ER_R operator to the input having more than one ER. The mutation operators of ACTesting are shown in Fig. 5.
Text Synthesizing. After completing the mutation operations based on the adaptability density-guided constraint, ACTesting constructs the synthesized text based on the mutated ER pools. To ensure the fluency of synthesized sentences and the integrity of semantic information, we utilize BERT (Devlin et al., 2019) to enhance the details of the sentences. At this point, we obtain paired text inputs.
Image object and relationship detection. The paired text inputs are sent to the tested T2I software respectively for generating the output images. To address the challenge of cross-modal semantic matching, we employ the object detection model to identify entities in images and use the scene graph model to detect relationships between entities in images during the implementation process. In the end, we forward the detected results to the next component.
Testing and Analysis. Based on the MR and mutation operators introduced in Sec 3.3 and 3.4, ACTesting computes the test results. We introduce the error-detection ratios and miss-detection ratios for entities and relationships to demonstrate the testing efficiency, detailed in Sec 4.3.
In summary, ACTesting encompasses a comprehensive process of new input text generation, image generation by T2I software and test result computation, where each module plays a crucial role in the overall effectiveness of the testing.
3.2. Adaptability Density-guided Constraint